Project

General

Profile

Actions

MOM update Nov 28, 2017

MOM update Oct 10, 2017

  • In the last week
    • Power outage & water works Mon-Wed
    • Overnight Tuesday 21/11:
      • Field-off, completed straight track data collection
    • Overnight Wednesday 22/11:
      • Alignment & tof calibration runs
    • Thursday, 23/11
      • SC magnets ramp up to ramped to 2017-02-2 by ~1500
      • 200 MeV running, solenoid mode
    • Friday 24/11
      • ~3pm, established 2017-02-6
      • 2T solenoid mode
      • 140 MeV runs + momentum scan
    • Saturday 25/11:
      • Tracker vacuum pump on Cryo 1 failed ~noon
      • Mark Tucker found a spare and put it on FC and moved the FC pump to Cryo 1.
      • Took calibration data after & continued with momentum scan at 2017-02-6
      • ISIS down ~1500-2100
      • Scaled down momentum scan data requirement. Completed overnight.
    • Sunday: ~2pm ramped to 2017-02-1, ran overnight
    • Monday 0600 Empty-LH2 data-taking ended
  • Issues:
    • Target lasers creeping down. Tweaked up power on B by ~5% on Saturday
    • Tracker Cryo1 vacuum pump failed -- caused temperatures to start creeping up a little [ second stage got up to ~7.25K ]
      • CM has spares for vac pump but not for turbo
    • Lost KL after power cut
      • Looks like HV crate had tripped off.
      • ALH didn't pick it up because the hv ioc is in a 'zombie' state (to be fixed)
      • Online recon summary canvas was missing KL (now fixed)
    • Hall Probes: one of the SSD hall probes is not updating on the b-field monitor. USB connection? IOC issue?
    • Need survey before LiH empty or latest before LiH goes in.

Updated by Rajaram, Durga about 6 years ago ยท 4 revisions